OpenStudy (anonymous):

Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing.

OpenStudy (e.mccormick):

Is that \(f(x)=20\cdot 1\cdot 3^x\)?

OpenStudy (anonymous):

yes

OpenStudy (e.mccormick):

Or is it supposed to be 1.3. Hard to tell on some of your dots if they are for multiplication or a decimal.

OpenStudy (anonymous):

20 times 1.3^x

OpenStudy (e.mccormick):

Now it says "here is the graph of the" then is that inverse? If so, not sure what it is an inverse of. Also, that is not the graph of \(f(x)=20\times 1.3^x\). You have (-5,10) marked, but it is close to (-5,5.39) that would be on the graph.

OpenStudy (anonymous):

3. Here your domain is # of years and the study lasted 15 years. Years will not be negative, This is the feedback from the teacher, Please help me I am confused

OpenStudy (e.mccormick):

Well, if it is a pure math formula, the domain is \((-\infty,\infty)\) and the range is \((0,\infty)\). But this is study data, so there is a restriction on the range.

OpenStudy (e.mccormick):

Sorry, restriction on the domain.

OpenStudy (anonymous):

So how I fix it in the graph and which one will be the domain, range and y intercept and how I graph that.

OpenStudy (e.mccormick):

The y intercept would not change. The restricted domain will cause a restricted range. In your chart near the top of the first page, you show 15 as giving 1024, which is what it rounds to. That is going to be your maximum y value.

OpenStudy (anonymous):

Thanks for your help I do not have medals but if I get one I will give it to you. I am very grateful for your help!!!!

OpenStudy (e.mccormick):

Just click where it says best response.

OpenStudy (e.mccormick):

See how this uses your chart and the calculation to make a graph? https://www.desmos.com/calculator/lcwlqyvuht

OpenStudy (anonymous):

THANK YOU I REALLY APPRECIATE IT!!!!!